# CP violation in two-meson Tau decays induced by heavy new physics

### v1

@@ 5.3 CP violating asymmetries in the $K^{\pm}\pi^{0}$ channel
The maximal $A_{FB}|_{K\pi}$ is shown by a solid blue line in figure 9, which was obtained for the following values of the Wilson coefficients: $\Re e[\epsilon_{S}^{s}]=2.3\times 10^{-2}$ , $\Im m[\epsilon_{S}^{s}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{s}]=1.8\times 10^{-2}$ and $\Im m[\epsilon_{T}^{s}]=4\times 10^{-6}$. The SM contribution is depicted by a black dashed line. The small difference between them seems to prevent any observation of NP in this observable in current or forthcoming experiments. In this case both curves are clearly non-vanishing, for most of the phase space, reflecting the effect of the kaon-pion mass difference on the angular distribution (as opposed to the channels discussed in sectiones 5.1 and 5.2, where this difference vanishes in the very good approximation of isospin symmetry).
@@
In the SM, the FoM $F(\hat{\epsilon}_{S}=0,\hat{\epsilon}_{T}=0)$ reads $0.45$. Varying the Wilson coefficients within their allowed ranges, we find that its maximum value is $0.64$, which is obtained for $\Re e[\epsilon_{S}^{s}]=2.3\times 10^{-2}$, $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-8\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=-4\times 10^{-6}$. For this maximal CP violation allowed, measurement shall be possible at present experiments. Although it may seem that differentiation from the SM result is fairly achievable, it would actually be challenging, according to Fig.  10, since it would require a quite accurate measurement in the $K^{*}(1410)$ region. Fig. 11 shows a three-dimensional density plot of the FoM $F$ for fixed $\Im m[\hat{\epsilon}_{T}^{s}]$, varying its real part and the complex number $\hat{\epsilon}_{S}^{s}$.
@@
For the $K\eta^{(\prime)}$ channels we obtain non-negligible $A_{CP}^{rate}$ ($\lesssim 10^{-5}$) and large deviations due to NP in $A_{FB}$ and the FoM $F$  [^15] These results will be reported elsewhere [106].. However, since these are inelastic channels for $K\pi^{0}$, it would be very demanding to control the SM prediction at the required level to claim NP using the asymmetries for these decay modes.
@@ 6 Conclusions
In this work we have examined –within the same EFT setting– the rate and forward-backward asymmetries for the other two-meson tau decay modes, focusing on the $(\pi/K)\pi^{0}$ and $KK_{S}$ cases. The most promising one turns out to be the di-Kaon channel, where measurements of $A_{CP}^{rate}$ with a $5\%$ precision would already be sensitive to NP contributions that are currently allowed. For these modes, also measurements of $A_{FB}$ and its integral in energy (that we quantify by our FoM $F$ introduced in eq. (33)) would easily distinguish NP permitted by present data. An additional and very timely interest of measuring $A_{CP}^{rate}$ for the $KK_{S}$ modes would be checking the result for $\pi K_{S}$, as they coincide unless ad hoc NP is invoked.
In the $\pi\pi^{0}$ channels $A_{FB}$ and $F$ would again be sensitive to NP in current and planned facilities. Disentangling NP from SM contributions would be however (much) more challenging in the $K\pi^{0}$ ($K\eta^{(\prime)}$) modes. In the first case because these deviations are mostly placed in the $K^{*}(1410)$ region and towards the end of the spectrum, where there are fewer events and data is less precise; and in the second because of the uncertainties on the $K\eta$ data ($\tau\to K\eta^{\prime}\nu_{\tau}$ decays have not been measured yet) and the NP effects being in the inelastic region of the coupled-channels $K(\pi^{0}/\eta/\eta^{\prime})$ problem, which is not so controlled.

### v2

@@ 5.3 CP violating asymmetries in the $K^{\pm}\pi^{0}$ channel
The maximal $A_{FB}|_{K\pi}$ is shown by a solid blue line in figure 9, which was obtained for the following values of the Wilson coefficients: $\Re e[\epsilon_{S}^{s}]=2.3\times 10^{-2}$ , $\Im m[\epsilon_{S}^{s}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{s}]=1.8\times 10^{-2}$ and $\Im m[\epsilon_{T}^{s}]=4\times 10^{-6}$. The SM contribution is depicted by a black dashed line. The small difference between them seems to prevent any observation of NP in this observable in current or forthcoming experiments. In this case both curves are clearly non-vanishing, for most of the phase space, reflecting the effect of the kaon-pion mass difference on the angular distribution (as opposed to the channels discussed in sections 5.1 and 5.2, where this difference vanishes in the very good approximation of isospin symmetry).
@@
In the SM, the FoM $F(\hat{\epsilon}_{S}=0,\hat{\epsilon}_{T}=0)$ reads $0.45$. Varying the Wilson coefficients within their allowed ranges, we find that its maximum value is $0.64$, which is obtained for $\Re e[\epsilon_{S}^{s}]=2.3\times 10^{-2}$, $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-8\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=-4\times 10^{-6}$. For this maximal CP violation allowed, measurement shall be possible at present experiments. Although it may seem that differentiation from the SM result is fairly achievable, it would actually be challenging, according to Figure 10, since it would require a quite accurate measurement in the $K^{*}(1410)$ region. Figure 11 shows a three-dimensional density plot of the FoM $F$ for fixed $\Im m[\hat{\epsilon}_{T}^{s}]$, varying its real part and the complex number $\hat{\epsilon}_{S}^{s}$.
@@
For the $K\eta^{(\prime)}$ channels we obtain small $A_{CP}^{rate}$ ($\lesssim 5\times 10^{-6}$) and non-negligible deviations due to NP in $A_{FB}$ and the FoM $F$  [^17] These results will be reported elsewhere DanielThesis .. However, since these are inelastic channels for $K\pi^{0}$, it would be very demanding to control the SM prediction at the required level to claim NP using the asymmetries for these decay modes.
@@ 6 Conclusions
In this work we have examined –within the same EFT setting– the rate and forward-backward asymmetries for the other two-meson tau decay modes, focusing on the $(\pi/K)\pi^{0}$ and $KK_{S}$ cases. The most promising one turns out to be the di-Kaon channel, where measurements of $A_{CP}^{rate}$ with a $5\%$ precision would already be sensitive to NP contributions that are currently allowed. For these modes, also measurements of $A_{FB}$ and its integral in energy (that we quantify by our FoM $F$ introduced in eq. (33)) would be able to distinguish NP permitted by present data. An additional and very timely interest of measuring $A_{CP}^{rate}$ for the $KK_{S}$ modes would be checking the result for $\pi K_{S}$, as they coincide (up to sign) unless ad hoc NP is invoked.
In the $\pi\pi^{0}$ channels $A_{FB}$ and $F$ would again be sensitive to NP in current and planned facilities, although difficult to measure. Disentangling NP from SM contributions would be however (much) more challenging in the $K\pi^{0}$ ($K\eta^{(\prime)}$) modes. In the first case because these deviations are mostly placed in the $K^{*}(1410)$ region and towards the end of the spectrum, where there are fewer events and data is less precise; and in the second because of the uncertainties on the $K\eta$ data ($\tau\to K\eta^{\prime}\nu_{\tau}$ decays have not been measured yet) and the NP effects being in the inelastic region of the coupled-channels $K(\pi^{0}/\eta/\eta^{\prime})$ problem, which is not so controlled.